### v2

# Enhancing Prohibited Item Detection through X-ray-Specific Augmentation and Contextual Feature Integration

Abstract
X-ray prohibited item detection faces challenges due to the long-tail distribution and unique characteristics of X-ray imaging. Traditional data augmentation strategies, such as copy-paste and mixup, are ineffective at improving the detection of rare items due to the complex interactions between overlapping objects. Furthermore, X-ray imaging removes easily distinguishable features like color and texture, making it difficult to differentiate between visually similar categories. To address these challenges, in this work, we propose the X-ray Imaging-driven Detection Network (XIDNet). Inspired by the unique characteristics of X-ray imaging, this network introduces two key innovations: a novel X-ray-specific augmentation strategy that generates more realistic training samples for rare items, thereby improving detection performance for categories with insufficient samples, and an contextual feature integration algorithm that captures the spatial and semantic interactions between objects and surroundings under X-ray imaging, enhancing the model’s ability to distinguish between similar categories. Extensive experimental results show that XIDNet effectively leverages X-ray imaging characteristics to significantly improve detection performance, outperforming popular SoTA methods by up to +17.2% in tail categories.
1 Introduction
As crowd density increases in public transportation hubs [2, 35], ensuring public safety through effective security inspections has become more critical than ever. X-ray scanners [40], widely deployed to inspect luggage and produce complex imagery, play a crucial role in these inspections. However, the accuracy of identifying prohibited items can be compromised due to the limitations of manual inspection, particularly under conditions of prolonged concentration, which may lead to missed detention of prohibited items and potential security breaches. Therefore, there is a pressing need for an accurate and automated detection to support these safety efforts. Advances in deep learning [18, 16], particularly CNNs, offer promising solutions by treating AI inspection as an object detection task [49, 48] in computer vision community [30].
Figure 1: Comparison of existing data augmentation methods for common and X-ray images. Traditional methods like copy-paste and mixup fail to capture X-ray-specific characteristics, leading to a noticeable gap between synthetic and real images.
Recent research has made significant progress in improving the detection accuracy of prohibited items in X-ray images. Several X-ray benchmarks, such as GDXray [22], SIXray [23], and OPIXray [39], have been introduced to provide valuable resources for this task. In real-world scenarios, the frequency of different categories varies, leading to a long-tail distribution in prohibited items detection data, as seen in datasets like HiXray [32]. On the model front, traditional CNN-based approaches [41, 9] have demonstrated promising results in X-ray object detection, leveraging advancements in deep learning and object recognition. Despite the challenges unique to X-ray imagery, such as noise and the absence of strong visual cues like color and texture, these methods have pushed the boundaries of detection performance, showing significant improvements in identifying a wide range of prohibited items.
However, despite these advancements, existing methods still face substantial limitations: (1) The unique imaging mechanism of X-rays often results in complex coupling between overlapping objects, making it challenging for traditional data augmentation strategies, such as copy-paste [11] and mixup [42], to effectively capture relationships (see Fig. 1). Moreover, these techniques may distort the environmental context of the X-ray original image, negatively impacting performance, particularly for tail categories. (2) Color filtering by X-ray imaging erases key distinguishing features, making it difficult to differentiate between visually similar objects (see Fig. 2). Different materials, such as metals, often appear nearly identical in X-ray scans, further complicating the task of accurate object identification.
To address these challenges, we propose the X-ray Imaging-driven Detection Network (XIDNet), which leverages the unique characteristics of X-ray imaging to improve detection performance for specialized categories of prohibited items. Specifically, we introduce two key innovations: a novel X-ray-specific augmentation strategy that utilizes Poisson blending to crop instances of rare categories from their original images and seamlessly integrate them into other images, generating realistic synthetic instances and improving detection performance for categories with few samples, and an environment-aware feature enhancement algorithm that captures the spatial and semantic interactions between objects and their surroundings in X-ray images, enhancing the model’s ability to distinguish between visually similar objects by leveraging contextual relationships between target items and their environment. Together, these innovations provide a comprehensive solution to the unique challenges of X-ray security inspection.
Figure 2: Preliminary experimental visualization. The left side shows the ground truth, while the right displays the predictions. We observe the difficulty in distinguishing between categories due to X-ray filtering. For example, when a hammer is present, its components are usually misclassified as a knife.
Extensive experimental evaluation demonstrates that XIDNet significantly outperforms existing SoTA methods, particularly in detecting these categories with insufficient samples. In summary, this work provides a comprehensive solution to the challenges posed by imbalanced data distributions in X-ray security screening, ultimately contributing to more reliable automated security systems. The main contributions are summarized as follows:
  • We are the first to utilize Poisson blending in X-ray-specific augmentation to generate more realistic X-ray samples, effectively addressing class imbalance and enhancing detection for tail categories with limited samples.
  • To overcome the challenge of category distinction due to color filtering in X-ray imaging, we propose a novel model-level strategy that enhances object discriminability by integrating surrounding contextual information.
  • Extensive experiments show that XIDNet outperforms existing methods by a substantial margin (+17.2%), providing a robust solution for accurate prohibited item detection in real-world X-ray security inspections.
Figure 3: The framework overview of XIDNet. The proposed XIDNet consists of two key components: X-ray-specific augmentation generates synthetic data by merging X-ray objects with backgrounds, closely mimicking real-world scenarios, and contextual feature integration, which strengthens proposal features using relational context from high-confidence proposals to improve overall performance.
2 Related Work
2.1 X-ray Prohibited Items Detection
X-ray imaging offers powerful ability in many tasks such as medical image analysis [12, 3, 20] and security inspection [23, 13]. As a matter of fact, obtaining X-ray images is difficult, so rare studies touch security inspection in computer vision due to the lack of specialized high-quality datasets.
Recently, multiple datasets have been made available for research, including SIXray [23], OPIXray [39], HiXray [32], EDS [33], X-ray FSOD [34], PIXray [21], PIDray [36], and CLCXray [46]. These datasets facilitate the evaluation of tasks such as image classification, object detection, cross-domain detection, and few-shot detection. However, a limitation of these datasets is that they only offer X-ray scans from a single viewpoint, which does not align with the dual-view X-ray scanning systems commonly used in real-world security applications. Although DBF [14] and MV-Xray [28] provide multi-view X-ray images, they are not publicly available due to privacy restrictions, which limits their use in the literature. GDXray [22] also includes multi-view X-ray images but focuses on individual objects rather than baggage, so it is considered a single-view dataset for the purposes of this study. Additionally, other X-ray datasets, such as those utilized in medical imaging [27], [38], differ significantly from security screening datasets, as objects in medical X-ray images typically do not overlap.
2.2 Long-tail Vision Tasks
Long-tail visual recognition addresses the imbalance between common (head) and rare (tail) classes in real-world datasets. Several approaches have been proposed to handle the challenge of learning effective representations for tail classes. Class rebalancing strategies like re-sampling and re-weighting aim to mitigate bias towards head classes. For example, [6] introduced an effective number of samples to create class-balanced loss weights, improving performance for underrepresented classes. Data augmentation techniques, such as [4], generate synthetic data to boost tail class instances and balance the class distribution. However, these methods may lead to overfitting or fail to capture the full diversity of tail classes.
Recent work, such as Logit Normalization (LogN) by [47], self-calibrates classified logits, similar to batch normalization, to address long-tail recognition. [45] proposed ROG, a method that reconciles object-level and global-level objectives in long-tail detection. Meta-learning and contrastive learning are also explored for improving feature learning in long-tail tasks. [15] introduced Decoupled Training, which first learns a feature extractor using cross-entropy loss and then fine-tunes it with a rebalanced classifier. [1] proposed LDAM Loss, which adjusts decision boundaries based on class frequency, increasing margins for tail classes. Additionally, contrastive learning methods like [31] align feature distributions to ensure consistency between head and tail classes, more robust to imbalance.
3 Method
In this section, we provide a detailed description of each component of the proposed XIDNet. We begin by introducing the overall framework in Sec. 3.1. Next, Sec. 3.2 covers the X-ray-specific augmentation, while Sec. 3.3 discusses the contextual feature integration. Finally, Sec. 3.4 outlines the complete network training process.
3.1 Framework Overview
Fig. 3 illustrates the proposed XIDNet framework, consisting of two main components: X-ray-specific augmentation and contextual feature integration.
The X-ray-specific augmentation module addresses class imbalance, especially between head and tail categories. Instances from tail classes are cropped from their original images and treated as foreground objects. These cropped instances are then seamlessly inserted into other images using Poisson image editing, ensuring natural integration with consistent lighting and background blending. This augmentation method enhances the representation of underrepresented classes, balancing the dataset and improving the model’s ability to generalize across rare categories.
The contextual feature integration module focuses on contextual relationships between objects. By utilizing spatial and semantic correlations between the target object and surrounding items, the model can better interpret complex scenes. For example, the presence of certain objects nearby may aid in accurate classification, especially when prohibited items are often surrounded by related items. This approach reduces misclassification by leveraging spatial dependencies, improving robustness of performance in cluttered or occluded object detection under X-ray imaging.
3.2 X-ray-Specific Augmentation
Traditional data augmentation techniques like Copy-paste [10] and Mixup [42] struggle with X-ray images due to material-dependent color and texture variations. Non-metallic objects are semi-transparent or orange, while metallic ones appear darker, complicating object overlap. As shown in Fig. 1, copy-paste causes significant discrepancies, while mixup disrupts the image context. To address this, we propose an X-ray-specific augmentation strategy that maintains contextual integrity while smoothly integrating new objects, as illustrated in Fig. 3 for tail classes.
Consider a dataset $D=\{(I_{k},B_{k},C_{k})\}_{k=1}^{K}$, where each image $I_{k}$ contains bounding boxes $B_{k}=\{b_{1}^{k},b_{2}^{k},\dots,b_{n_{k}}^{k}\}$ and corresponding class labels $C_{k}=\{c_{1}^{k},c_{2}^{k},\dots,c_{n_{k}}^{k}\}$. In a typical long-tail distribution, the classes are divided into $C=C_{head}\cup C_{tail}$, where $C_{head}$ represents frequently occurring classes, and $C_{tail}$ represents the rarely occurring tail classes. The instances of tail classes, $N_{c}$, are much fewer than those of head classes, $N_{c^{\prime}}$, i.e., $N_{c}\ll N_{c^{\prime}}$, which limits the model’s ability to effectively generalize. For each tail class $c\in C_{tail}$, we perform the following steps:
Instance Cropping. For each image $I_{k}$ containing an instance of class $c$, the bounding box $b_{i}^{k}\in B_{k}$ corresponding to class $c$ is cropped to generate a sub-image $I_{c,i}=I_{k}[b_{i}^{k}]$. The cropped instances form the set $S_{c}=\{I_{c,i}\}$. Since the existing bounding boxes often include background elements, we manually annotate segmentation masks for precise cropping, ensuring only foreground instances are used (further details are in the supplementary materials).
Position Selection. After cropping the instances, we place them in positions corresponding to existing objects in the image, using the object labels to determine appropriate locations. This prevents placing objects in blank areas and ensures realistic overlap, as shown in Fig. 1. This method, combined with the smooth fusion process, generates varied and complex overlapping cases that improve model performance on challenging scenarios.
Poisson Fusion. We use Poisson fusion [24] for its ability to accurately model the interaction between X-ray objects and their backgrounds. Poisson fusion preserves important X-ray details, such as texture and contrast, by ensuring smooth integration of the foreground object with the background. Each cropped instance $I_{c,i}\in S_{c}$ is inserted into a new background image $I_{bg}$ using Poisson fusion to create a seamless composite image $I_{new}$, represented as:
$I_{new}=\text{PoissonFusion}(I_{bg},I_{c,i},b_{new}),$ (1)
where $b_{new}$ denotes new bounding box of inserted object.
Augmented Dataset. The newly generated images $I_{new}$, along with bounding boxes and labels, are added to the training dataset, resulting in an augmented dataset $D^{\prime}$:
$D^{\prime}=D\cup\{(I_{new},B_{new},C_{new})\},$ (2)
where $B_{new}$, $C_{new}$ are augmented bounding boxes, labels.
By applying the X-ray-specific augmentation to tail class instances, we increase the effective number of training samples, reducing class imbalance and improving the model’s ability to generalize to tail classes during training.
3.3 Contextual Feature Integration
In this module, we enhance the detection process by leveraging the contextual relationships between objects and surroundings in the image. The first step is to select the top $k$ proposals based on their confidence scores. Let the set of initial proposals be $\{p_{1},p_{2},\dots,p_{N}\}$, with corresponding confidence scores $\{s_{1},s_{2},\dots,s_{N}\}$. We then select the top $k$ proposals, denoted by $P_{\text{top-k}}=\{p_{i_{1}},p_{i_{2}},\dots,p_{i_{k}}\}$, where the confidence scores satisfy $s_{i_{1}}\geq s_{i_{2}}\geq\dots\geq s_{i_{k}}$. These proposals are assumed to contain the most likely prohibited items and will serve as the basis for further enhancement.
For each of these top-$k$ proposals, we extract their feature vectors, denoted as $f_{i_{j}}$ for $j=1,2,\dots,k$. To capture the relationships among these top proposals, we concatenate their feature vectors into a single vector:
$F_{\text{concat}}=\text{Concat}(f_{i_{1}},f_{i_{2}},\dots,f_{i_{k}})$ (3)
This concatenated feature vector $F_{\text{concat}}$ contains information about the dependencies between the top-$k$ proposals, reflecting how they interact in the context of the image. We then apply a fully connected layer to process $F_{\text{concat}}$ and produce a relational fusion feature, $F_{\text{fusion}}$, as follows:
$F_{\text{fusion}}=\sigma(\mathbf{W}_{1}F_{\text{concat}}+\mathbf{b}_{1}),$ (4)
where $\sigma$ is an activation function, such as ReLU, and $\mathbf{W}_{1}$ and $\mathbf{b}_{1}$ are the weights and bias of the fully connected layer. The relational feature $F_{\text{fusion}}$ captures the higher-order interactions between the selected proposals, encoding the contextual relationships that are crucial for distinguishing between visually similar categories.
To further integrate this relational information, we concatenate the fused relational feature $F_{\text{fusion}}$ with each of the original proposal features $f_{i_{j}}$, forming an augmented feature vector for each proposal. We then apply another fully connected layer to each augmented feature vector to produce the final updated feature vector $f_{i_{j}}^{\text{new}}$. This process can be formulated as follows:
$f_{i_{j}}^{\text{aug}}=\text{Concat}(f_{i_{j}},F_{\text{fusion}}),$ (5)
$f_{i_{j}}^{\text{new}}=\sigma(\mathbf{W}_{2}f_{i_{j}}^{\text{aug}}+\mathbf{b}_{2}),$ (6)
where $\mathbf{W}_{2}$ and $\mathbf{b}_{2}$ are the weights and bias of the second fully connected layer. This updated feature vector $f_{i_{j}}^{\text{new}}$ incorporates both the original features of the proposals and the relational information from other proposals, enabling the model to make more informed classification decisions.
Finally, the updated features $f_{i_{j}}^{\text{new}}$ replace the original proposal features in the classification and bounding box regression heads. The new classification logits are computed from these enhanced features:
$\text{Logits}_{\text{new}}=\text{ClassificationHead}(f_{i_{j}}^{\text{new}})$ (7)
3.4 Overall Training Process
Algorithm 1 outlines the entire training procedure for XIDNet, highlighting the key stages of the X-ray-Specific augmentation and Contextual Feature Integration, along with the overall training process. The X-ray-Specific augmentation module involves cropping instances and applying Poisson fusion to generate synthetic data for tail classes. The Contextual Feature Integration module improves feature representations of proposals by leveraging relational context from high-confidence proposals. Finally, the detection model is trained using the augmented dataset and context-enhanced features. By integrating these two modules, XIDNet addresses the long-tail distribution problem in X-ray prohibited items detection, delivering superior performance, especially for tail categories.
Input: Input images $I_{k}$, bounding boxes $B_{k}$, class labels $C_{k}$, for all $k\in[1,K]$.
Output: Trained model $\mathbf{M}$.
X-ray-Specific Augmentation:
for each tail class $c\in C_{tail}$ do
       for each instance $b_{i}^{k}\in B_{k}$ of class $c$ do
             Crop $I_{c,i}=I_{k}[b_{i}^{k}]$;
             Generate $I_{new}=\text{PoissonFusion}(I_{bg},I_{c,i},b_{new})$;
             Add $(I_{new},B_{new},C_{new})$ to dataset $D^{\prime}$;
Contextual Feature Integration:
Sample Batches from the $D^{\prime}$;
Generate proposals $\{p_{1},p_{2},\dots,p_{N}\}$ using RPN, with scores $\{s_{1},s_{2},\dots,s_{N}\}$;
Select $P_{\text{top-k}}=\{p_{i_{1}},p_{i_{2}},\dots,p_{i_{k}}\}$ based on scores;
Concatenate $F_{\text{concat}}=\text{Concat}(f_{i_{1}},f_{i_{2}},\dots,f_{i_{k}})$;
Compute fusion $F_{\text{fusion}}=\sigma(\mathbf{W}_{1}F_{\text{concat}}+\mathbf{b}_{1})$;
for each feature vector $f_{i_{j}}$ in $P_{\text{top-k}}$ do
       Compute $f_{i_{j}}^{\text{aug}}=\text{Concat}(f_{i_{j}},F_{\text{fusion}})$;
       Update feature $f_{i_{j}}^{\text{new}}=\sigma(\mathbf{W}_{2}f_{i_{j}}^{\text{aug}}+\mathbf{b}_{2})$;
Replace original logits with $\text{Logits}_{\text{new}}=\text{ClassificationHead}(f_{i_{j}}^{\text{new}})$;
Perform backpropagation and update parameters;
Output the trained model $\mathbf{M}$.
Algorithm 1 Training Procedure of XIDNet
4 Experiments
We provide a comprehensive evaluation of XIDNet. Sec. 4.1 outlines the experimental setup. Sec. 4.2 evaluates the effectiveness and generalizability of integrating it into various baseline models. Sec. 4.3 compares XIDNet with SoTA long-tail detection techniques, while Sec. 4.5 contrasts it with popular X-ray detection methods. Sec. 4.6 presents ablation studies to assess the contributions of each module, as well as the impact of the parameter $k$. Finally, Sec. 4.7 includes experiments on manual segmentation scalability and generalization to transformer-based models.
Method HiXray Dataset PIDray Dataset
AP50 PO1 PO2 WA LA MP TA CO NL (Tail) AP50 BA PL HA PO SC WR GU BU SP HA KN LI
F-RCNN [25] 84.2 95.0 93.4 91.9 98.0 97.0 95.7 71.0 32.0 76.0 83.7 96.0 88.5 60.7 91.2 96.7 47.6 71.9 67.3 98.7 42.1 67.8
+ ours 85.2 96.1 94.4 92.4 98.0 97.0 94.1 68.0 41.7↑9.7% 77.6 87.7 96.4 90.0 63.2 92.5 97.0 48.2 72.6 72.8 98.7 42.8 69.6
S-RCNN[29] 80.8 94.2 92.4 90.3 98.2 97.0 95.1 63.6 15.1 71.8 84.4 95.0 86.7 62.3 86.4 96.8 37.1 66.0 48.6 98.5 33.9 65.7
+ ours 83.2 94.8 93.6 91.2 97.9 97.3 94.5 66.5 29.7↑14.6% 72.4 85.0 94.5 83.3 63.5 86.3 96.3 38.5 66.8 56.0 98.6 34.3 66.5
RetinaNet[26] 79.0 94.0 93.2 90.8 98.1 97.3 95.4 61.8 1.7 69.0 77.6 93.1 82.8 48.9 83.7 91.6 38.4 64.2 58.2 98.0 28.0 63.5
[…]
CenterNet[8] 82.4 95.6 94.3 90.7 98.1 97.8 95.3 73.1 14.6 74.0 86.5 96.8 90.3 56.9 90.6 95.6 35.3 68.7 67.9 98.5 35.3 65.7
+ ours 83.7 95.5 94.5 90.9 98.0 97.9 95.4 73.2 24.3↑9.7% 74.3 87.0 96.9 91.4 54.7 90.3 97.0 36.9 72.8 66.7 98.4 33.3 65.8
ATSS[44] 83.5 95.2 93.6 91.8 98.2 97.4 95.4 68.0 28.3 74.6 84.1 96.3 91.2 56.5 90.9 96.4 39.8 70.6 71.4 98.5 33.0 66.7
+ ours 85.6 95.7 94.2 93.1 98.1 97.4 95.4 70.6 40.7↑12.4% 76.2 87.6 96.9 90.0 58.2 90.0 96.6 49.3 72.1 74.7 98.0 32.8 67.8
Table 1: Comparison of detection performance between baselines and XIDNet-enhanced (+ours) versions on the HiXray and PIDray datasets. The proposed XIDNet consistently improves AP50 of the overall categories, especially for these tail ones in HiXray dataset.
4.1 Experimental Setup
Tasks and Datasets. To ensure fair comparisons with previous detection methods, we evaluate our model on two large-scale X-ray datasets, HiXray [32] and PIDray [36]. HiXray is derived from real-world scenarios and features a distinctly long-tail distribution, making it ideal for evaluating model performance on naturally imbalanced data. Although PIDray samples are synthetically generated and balanced, this dataset offers a comprehensive evaluation framework for X-ray imaging tasks and is suitable for testing the classification capabilities of the proposed XIDNet.
Evaluation Metrics. Following standard object detection practices, we use Average Precision (AP) to assess the model’s category-specific performance and mean Average Precision (mAP) for overall effectiveness. AP50 is calculated by averaging across ten Intersection over Union (IoU) thresholds 0.5, while mAP represents the mean of these AP values across all evaluated categories, providing a robust measure of the model’s overall accuracy on detecting items.
Implementation Details. All the experiments were conducted using the mmdetection open-source toolbox [^1] https://github.com/open-mmlab/mmdetection [5] to ensure consistency and comparability. The parameter settings for all models followed the default configurations of mmdetection. Experiments were conducted on four NVIDIA GeForce RTX 4090 GPUs.
4.2 Comparison with Various Detection Baselines
To determine the performance and generalizability of XIDNet, we integrated it with multiple popular object detection frameworks, including F-RCNN (Faster RCNN [25]), S-RCNN (Sparse RCNN [29]), RetinaNet [26], CenterNet [8], and ATSS [44]. The results are summarized in Tab. 1.
Performance on HiXray Dataset. Tab. 1 shows that XIDNet consistently improved detection accuracy across all baseline models on the HiXray dataset. For Faster RCNN, the baseline achieved an $\text{AP}_{50}$ of 84.2%, whereas integrating XIDNet increased it to 85.2%. The most significant gains were observed in tail classes like class NL, which saw an improvement from 32.0% to 41.7%, an increase of 9.7%. This demonstrates XIDNet’s strength in boosting the performance of tail classes. For ATSS, XIDNet provided an $\text{AP}_{50}$ of 85.6%, compared to the baseline’s 83.5%. Specific improvements included: (1) Class NL: Improved from 28.3% to 40.7%, which is crucial given that NL is a challenging tail class with limited representation. (2) Class PO2: AP increased from 93.6% to 94.2%, which, although modest, shows improved consistency in handling frequent classes. Overall, the AP50 for ATSS improved by 2.1%, demonstrating that XIDNet’s dual-level augmentations effectively balance the performance between head and tail classes.
Performance on PIDray Dataset. The XIDNet also demonstrated strong performance gains on the PIDray dataset. Faster RCNN, for instance, saw an increase in AP50 from 76.0% to 77.6%. Notable improvements include: (1) The AP of class GU improved from 47.6% to 48.2%. This improvement, while relatively small, demonstrates XIDNet’s ability to provide more robust features for objects that may be occluded or embedded within a complex context. Guns, in particular, often have parts that overlap with other objects, and the improvement suggests better discrimination for such features. (2) The AP of class SP rose from 67.3% to 72.8%. This significant increase indicates that XIDNet’s augmentations made the model better at recognizing screwdrivers even when partially occluded or in cluttered environments, which is a common scenario in X-ray imagery. Across all baseline models, XIDNet’s integration led to increased detection performance for tail classes, highlighting the robustness of both data-level and model-level augmentation strategies. For instance, using S-RCNN with XIDNet on the PIDray dataset resulted in an increase in AP for GU from 37.1% to 38.5%, which is particularly notable given the scarcity of instances for this class. This shows XIDNet’s capability in enhancing detection for categories with insufficient training samples.
4.3 Comparison with Long-Tail Detection Methods
We compared XIDNet with several sota long-tail task methods. The methods considered in this comparison are LogN [47], ROG [45], Seasaw [37], GHM [19], and Focal [26]. These methods are popular for addressing the challenges posed by class imbalances, especially for the tail category.
XIDNet outperforms all the methods in terms of overall detection performance, achieving an AP50 score of 85.2, surpassing ROG (84.7) and Seasaw (83.8). The proposed framework integrates both X-ray-specific augmentation and contextual feature integration, making it particularly effective in overcoming the challenges posed by long-tail distributions. These innovations allow XIDNet to not only perform well on head categories but also provide significant improvements for tail categories, where most other methods are struggle to achieve the balance.
Method AP50 AP50 across various categories
PO1 PO2 WA LA MP TA CO NL
LogN [47] 83.1 95.2 93.6 92.0 98.0 96.7 95.1 62.8 31.3
ROG [45] 84.7 95.8 94.0 93.1 98.2 97.5 96.3 67.3 35.2
Seasaw [37] 83.8 95.4 94.3 92.3 98.1 97.0 88.3 69.5 35.5
GHM [19] 80.4 93.8 90.4 88.7 97.3 96.5 92.4 53.7 30.1
Focal [26] 82.2 94.7 93.2 89.5 97.5 97.6 91.5 63.0 30.2
XIDNet (ours) 85.2 96.1 94.4 92.4 98.0 97.0 94.1 68.0 41.7
Table 2: Comparison with popular long-tail detection methods on the HiXray dataset. Results highlight the advantage of XIDNet in addressing data imbalance, particularly for rare prohibited items.
Focusing on head categories, XIDNet maintains strong performance, achieving 96.1 in PO1 and 94.4 in PO2. This is higher than ROG, which scores 95.8 and 94.0, and Seasaw, which achieves 95.4 and 94.3. This indicates that XIDNet is on par with state-of-the-art methods in handling the majority classes. However, its true strength lies in the detection of tail categories, particularly in the detection of NL category, where XIDNet scores 41.7, a substantial improvement over ROG (35.2) and Seasaw (35.5). This highlights XIDNet’s ability to address the severe class imbalance typically encountered in X-ray prohibited item detection.
The performance of XIDNet in tail category detection is significantly better than conventional methods, particularly GHM (30.1) and Focal (30.2), which have a much lower AP50 in the tail class. XIDNet’s ability to generate synthetic data for tail classes through Poisson blending and its integration of contextual information from surrounding objects in the image contribute to its superior performance. This combined approach leads to enhanced feature representations for tail items, allowing the model to better handle occlusions and complex scenes. Therefore, XIDNet not only excels in detecting common prohibited items but also proves to be highly effective for rare items, making it a robust solution for X-ray screening in security applications.
4.4 Comparison with Data Augmentation Methods
In this section, we compare the performance of XIDNet with several popular data augmentation methods, including Copy-paste [11], Mixup [42], and Cutout [7]. These methods are widely used for enhancing the performance of object detection models, especially in the context of imbalanced datasets. The results are illustrated in Tab. 3.
The Copy-paste augmentation method achieves an impressive average AP50 of 84.4, with particularly strong performance across the head categories (PO1, PO2) and moderate performance for tail categories. However, as expected, its performance drops significantly in tail categories such as NL (26.4), highlighting the challenge of handling rare classes with this method. For fairness, we also used manually optimized instances in the Copy-paste experiment. Mixup yields an AP50 of 80.8, which is lower than Copy-paste. It performs well in head categories, particularly PO1 (94.4) and PO2 (92.8), but similarly struggles with tail categories, with a relatively low performance for NL (16.3). Cutout provides an AP50 of 83.9, ranking between Copy-paste and Mixup in overall performance. Its performance in head categories is competitive (e.g., PO1: 95.5, PO2: 94.0), but, like other augmentation methods, it experiences a significant drop in performance for tail category (NL: 25.2).
All data augmentation methods perform well in head classes (PO1, PO2) and achieve high AP50 scores for these categories. Copy-paste leads with the highest overall AP50 and performs particularly well across multiple head categories. The most significant challenge for these augmentation methods is improving the performance of tail categories (e.g., NL). Copy-paste achieves the best performance in this category compared to Mixup and Cutout, but it still falls short of what is needed for a robust tail-class detection.
Method AP50 AP50 across various categories
PO1 PO2 WA LA MP TA CO NL
Copy-paste [11] 84.4 96.3 95.7 94.1 98.0 98.2 96.4 69.7 26.4
Mixup [42] 80.8 94.4 92.8 89.3 97.9 96.9 94.2 64.8 16.3
Cutout [7] 83.9 95.5 93.9 91.6 98.2 97.5 96.2 73.2 25.2
DOAM[39] 82.2 95.9 94.7 93.7 98.1 98.1 95.8 65.0 16.1
LIM [32] 83.2 96.1 95.1 93.9 98.2 98.3 96.4 65.8 21.3
XIDNet (ours) 83.5 94.5 93.3 92.0 98.1 97.2 94.5 67.7 31.0
Table 3: Comparison with popular augmentation methods and X-ray-specific detection methods on HiXray (base model: ATSS).
4.5 Comparison with X-ray Detection Methods
In this section, we compare the performance of XIDNet with two popular X-ray prohibited items detection methods: DOAM [39] and LIM [32], which are specifically designed for X-ray images. The results are illustrated in Tab. 3.
The DOAM method achieves an AP50 of 82.2, showing strong performance for head categories (PO1: 95.9, PO2: 94.7). However, it suffers from significant degradation in performance for tail categories, especially NL (16.1), highlighting the difficulty of handling the categories with insufficient training samples. Moverever, LIM outperforms DOAM, with an AP50 of 83.2. It performs similarly in head categories, achieving strong results across PO1 (96.1), PO2 (95.1), and others, but still faces challenges with tail categories, such as NL (21.3), though its performance in tail classes is slightly better than the DOAM method.
Both DOAM and LIM exhibit strong performance in detecting head categories, with LIM achieving slightly better results overall. However, neither method achieves the robustness needed to handle tail classes effectively. XIDNet outperforms both DOAM and LIM in tail class performance, with a significantly higher AP50 of 31.0 in the NL category. This demonstrates the effectiveness of XIDNet’s data augmentation strategies and the contextual feature integration module, which are designed to improve detection accuracy in cluttered or occluded scenes, especially for the categories with insufficient training samples.
4.6 Ablation Studies
We conducted ablation studies to determine the contributions of data-level and model-level modules individually, as well as the impact of hyperparameter $k$ of the model-level.
4.6.1 Effectiveness of Each Module
The ablation results of each module are provided in Tab. 5. We analyzed the effects of enabling data-level augmentation, model-level augmentation, and both.
Setting AP50 AP50 across various categories
D M PO1 PO2 WA LA MP TA CO NL
84.2 95.0 93.4 91.9 98.0 97.0 95.7 71.0 32.0
84.6 95.3 94.5 92.5 98.0 97.0 94.8 68.0 36.8
85.0 95.3 93.9 92.3 97.9 97.7 94.3 67.4 41.2
85.2 96.1 94.4 92.4 98.0 97.0 94.1 68.0 41.7
Table 4: Ablation study on individual contributions of data-level (D) and model-level (M) modules. The table demonstrates how each component contributes to the overall detection improvement.
As illustrated in Tab. 5, Baseline with no augmentation achieved an AP50 of 84.2%. This served as our control, providing a point of comparison to assess the individual and combined contributions of each module. Data-Level Augmentation only improved to 85.0%. The class-specific performance showed notable gains for tail classes, such as Class NL which increased from 32.0% to 41.2%. This increase illustrates how data-level augmentation directly addresses data scarcity issues by providing more diverse training samples for underrepresented classes. Model-Level Strategy only resulted in an AP50 of 84.6%. The smaller overall improvement compared to data-level augmentation suggests that model-level augmentation is more complementary rather than a complete solution by itself. Combined Augmentation with full XIDNet achieved the best performance, with an AP50 of 85.2%. Notably, Class PO1 improved from 95.0% to 96.1%, and Class PO2 increased from 93.4% to 94.4%. These improvements confirm that both augmentation strategies are highly effective when combined, each addressing different aspects of the long-tail distribution problem.
4.6.2 Impact of Hyperparameters
To evaluate the sensitivity of the augmentation strength parameter $k$ on model-level augmentation, we experimented with $k$ values of 2, 4, 8, and 16 on RetinaNet [26], as shown in Fig. 4. $k$ equals 2 achieved an AP50 of 81.3%. The lower value of $k$ limits the amount of contextual information available for the augmentation, which results in less comprehensive relational learning. $k$ equals 4 yielded the highest AP50 of 81.8%, demonstrating that moderate incorporation of contextual features strikes the best balance between providing meaningful relationships without overwhelming the model with noise. $k$ equals 8 resulted in a slight decline to AP50 of 81.6%. Here, the increase in contextual information started to introduce noise, which diluted the effectiveness of the relational augmentation, indicating that too much contextual data can make feature learning less focused. $k$ equals 16 further decreased the performance to AP50 of 81.0%, with performance for tail classes like Class NL dropping to 14.6%. This suggests that excessive augmentation leads to model confusion, especially for tail classes in real-world scenario, as it becomes harder for the model to distinguish between relevant and irrelevant contextual cues of surroundings.
Figure 4: Performance impact of the $k$ on the model-level module. Results show optimal performance when $k$ = 4.
4.7 More Analysis and Discussion
This section discusses the scalability of manual segmentation and the generalization of to transformer-based models.
4.7.1 Manual segmentation scalability
In our initial experiments, we tested SAM [17] for automatic segmentation, but the performance was significantly lower than when manually annotating the objects. This discrepancy is particularly noticeable in terms of accuracy, as SAM [17] struggled with distinguishing some fine-grained details, leading to less precise segmentations (examples are in Fig. 5). However, the manual annotation of 78 instances, as shown in the supplementary material, took less than an hour, which is far less than the time required for a single training epoch. Given this, the manual segmentation process presents an efficient trade-off between annotation time and the quality of augmented data, making it a practical solution for the current scope of our work.
Figure 5: SAM automatic annotation (left: original, right: SAM).
4.7.2 Generalization on Transformer-based Models
We further evaluated the performance of our method by incorporating it with a transformer-based model, DINO [43]. As shown in Tab 5, our augmentation strategy leads to a noticeable improvement in detection performance. Specifically, the baseline DINO model achieves an AP50 of 84.4, whereas the integration of our technique boosts this to 86.3, with a significant increase in tail categories (NL) from 26.4 to 37.7. This demonstrates that our approach is effective not only for traditional convolutional models but also generalizes well to transformer-based models, improving overall performance, particularly for underrepresented categories.
Method AP50 AP50 across various categories
PO1 PO2 WA LA MP TA CO NL
DINO [43] 84.4 96.3 95.7 94.1 98.0 98.2 96.4 69.7 26.4
+ours 86.3 96.2 96.1 93.7 98.3 98.2 96.1 74.0 37.7
Table 5: Comparison with the transformer-based model, DINO.
5 Conclusion
In conclusion, the proposed XIDNet addresses critical challenges in X-ray prohibited item detection, particularly the issues stemming from the long-tail distribution of item appearances and the inherent complexities of X-ray imagery. By leveraging a dual-level boost framework, combining data augmentation via Poisson blending and a context-aware feature enhancement module, XIDNet significantly improves the detection of rare, infrequent items while maintaining high accuracy for common ones. Extensive experiments demonstrate that XIDNet outperforms state-of-the-art approaches, especially in enhancing detection for underrepresented classes. The contribution of this paper marks a significant advancement towards automated and robust security inspection systems, ensuring enhanced public safety in real-world environments like public transportation hubs.